# Scattering and induced false vacuum decay in the two-dimensional quantum Ising model

### v2

@@ Abstract
Spin chains are the go-to models for scattering simulations, as they can be efficiently simulated with tensor networks [17, 18, 19, 20, 21]. More importantly, they also emerge as effective descriptions of more complex models and phenomena. A prominent example is confinement, a non-perturbative mechanism that binds quarks into hadrons [22]. Analogous phenomena exist in the Ising model in one [23, 24] and two dimensions [25, 26, 27, 28, 29], where excitations feel a confining potential that binds them into bound states, reminiscent of mesons. Simulating scattering in spin chains can deepen our understanding of confining theories [7, 10], and quantum field theory [2, 30, 31]. Understanding the nature of composite excitations is also important in strongly-correlated condensed matter, like high-temperature superconductors [32, 33].
@@ II From elastic to inelastic scattering
One feature that stands out is the spatial distribution of the products: they dominantly disperse in two directions. This is a consequence of the shape of the magnon dispersion. The probability amplitude for scattering into a final state is proportional to the density of its states at the given energy, given by $|\nabla_{\mathbf{k}}E|^{-1}$ for dispersion $E$. The magnon dispersion, Eq. (2), produces a density of states with divergences – known as van Hoove singularities – at $k_{x}=k_{y}=\pm\frac{\pi}{2}$ and $k_{x}=-k_{y}=\pm\frac{\pi}{2}$. This greatly increases the probability for scattering in these directions. Note that this only applies if the scattering generates a pair of magnons. This indicates that the process at $g/J\sim 1.5$ must be two magnons $\rightarrow$ intermediate excitation $\rightarrow$ two magnons. The effect fades away at larger $g/J$, hinting at a different inelastic process in that regime.
@@
The evolution of the correlations is shown in Fig. 2(e), and $g$-dependence of peak amplitudes in Fig. 2(f). At $g/J=1$, the single dominant peak is a signature of the elastic process. It splits into two with increasing $g$, and at $g/J\gtrsim 1.5$ we clearly find two separate contributions. The first still corresponds to the elastic process, while the second signals the delayed arrival of particles generated through an intermediate resonance. The second peak is accompanied by a peak in $c_{O}$. The peaks of $c_{A}$ and $c_{O}$ are not equal because the magnon dispersion is not exactly sinusoidal, and thus the scattering is not precisely at the van Hoove singularity. Note that the products are more concentrated in the off-axis direction for $g/J=1.5$. We expect that this is accompanied by a finite-size effect, as the products are not point-like particles, but rather extended objects on the lattice. The negative $c_{AO}$ that accompanies the overlap of $c_{A}$ and $c_{O}$ shows that the presence of particles on the two axis is anti-correlated. This confirms that these are two separate two-particle processes, and not a four-particle process.
@@ III Scattering in a false vacuum
Originating in cosmology and high-energy physics [44], the question has recently attracted attention in the context of metastability in quantum many-body systems. It was studied numerically [45, 30, 46] and experimentally [47, 48, 49] in spin chains in 1D. The microscopic mechanism behind the decay is probabilistic bubble nucleation [46]: the vacuum fluctuations induced through a quench are small bubbles of the true vacuum, where the system locally tunnels into the stable state. Assuming a spherical bubble with radius $r$, its energy consists of two contributions: a positive interface tension, and a negative bulk term coming from the energy difference between the two vacua [44]. In 2D one expects the energy $E(r)=A\cdot 2J\cdot r-B\cdot 2h\cdot r^{2}$, with model-dependent geometrical factors $A$ and $B$. The parabola acts as a potential barrier between the false vacuum at $r=0$ and the true vacuum at $r\rightarrow\infty$, with the maximal energy at the critical bubble size $r^{*}\propto J/h$ and $E(r^{*})\propto J^{2}/h$. If a bubble bigger than $r^{*}$ is created, it can expand by dissipating energy, and the system transitions into the stable manifold. The energy manifold is sketched in Fig. 3(a).
@@ Acknowledgements
LP thanks Lorenzo Maffi and Pietro Silvi for valuable discussions, and Marco Rigobello for insightful comments on the manuscript. The tensor network simulations were performed with the Quantum Tea library [53, 54]. The exact diagonalisation calculations were performed with QuSpin [55].

### v3

@@ Abstract
Spin chains are the go-to models for scattering simulations, as they can be efficiently simulated with tensor networks [17, 18, 19]. More importantly, they also emerge as effective descriptions of more complex models and phenomena. A prominent example is confinement, a non-perturbative mechanism that binds quarks into hadrons [20]. Analogous phenomena exist in the Ising model in one [21, 22] and two dimensions [23, 24, 25, 26, 27], where excitations feel a confining potential that binds them into bound states, reminiscent of mesons. Simulating scattering in spin chains can deepen our understanding of confining theories [7, 10], and quantum field theory [2, 28, 29]. Understanding the nature of composite excitations is also important in strongly-correlated condensed matter, like high-temperature superconductors [30, 31].
@@ II From elastic to inelastic scattering
One feature that stands out is the spatial distribution of the products: they dominantly disperse in two directions. This is a consequence of the shape of the magnon dispersion. The probability amplitude for scattering into a final state is proportional to the density of its states at the given energy, given by $|\nabla_{\mathbf{k}}E|^{-1}$ for dispersion $E$. The magnon dispersion, Eq. (2), produces a density of states with divergences – known as van Hove singularities – at $k_{x}=k_{y}=\pm\frac{\pi}{2}$ and $k_{x}=-k_{y}=\pm\frac{\pi}{2}$. This greatly increases the probability for scattering in these directions. Note that this only applies if the scattering generates a pair of magnons. This indicates that the process at $g/J\sim 1.5$ must be two magnons $\rightarrow$ intermediate excitation $\rightarrow$ two magnons. The effect fades away at larger $g/J$, hinting at a different inelastic process in that regime.
@@
The evolution of the correlations is shown in Fig. 2(e), and $g$-dependence of peak amplitudes in Fig. 2(f). At $g/J=1$, the single dominant peak is a signature of the elastic process. It splits into two with increasing $g$, and at $g/J\gtrsim 1.5$ we clearly find two separate contributions. The first still corresponds to the elastic process, while the second signals the delayed arrival of particles generated through an intermediate resonance. The second peak is accompanied by a peak in $c_{O}$. The peaks of $c_{A}$ and $c_{O}$ are not equal because the magnon dispersion is not exactly sinusoidal, and thus the scattering is not precisely at the van Hove singularity. Note that the products are more concentrated in the off-axis direction for $g/J=1.5$. We expect that this is accompanied by a finite-size effect, as the products are not point-like particles, but rather extended objects on the lattice. The negative $c_{AO}$ that accompanies the overlap of $c_{A}$ and $c_{O}$ shows that the presence of particles on the two axis is anti-correlated. This confirms that these are two separate two-particle processes, and not a four-particle process.
@@ III Scattering in a false vacuum
Originating in cosmology and high-energy physics [42], the question has recently attracted attention in the context of metastability in quantum many-body systems. It has been studied in 1D numerically [43, 44, 28, 45, 46] and experimentally [47, 48, 49, 50]. The microscopic mechanism behind the decay is probabilistic bubble nucleation [45]: the vacuum fluctuations induced through a quench are small bubbles of the true vacuum, where the system locally tunnels into the stable state. Assuming a spherical bubble with radius $r$, its energy consists of two contributions: a positive interface tension, and a negative bulk term coming from the energy difference between the two vacuua [42]. In 2D one expects the energy $E(r)=A\cdot 2J\cdot r-B\cdot 2h\cdot r^{2}$, with model-dependent geometrical factors $A$ and $B$. The parabola acts as a potential barrier between the false vacuum at $r=0$ and the true vacuum at $r\rightarrow\infty$, with the maximal energy at the critical bubble size $r^{*}\propto J/h$ and $E(r^{*})\propto J^{2}/h$. If a bubble bigger than $r^{*}$ is created, it can expand by dissipating energy, and the system transitions into the stable manifold. The energy manifold is sketched in Fig. 3(a).
@@ Acknowledgements
LP thanks Lorenzo Maffi and Pietro Silvi for valuable discussions, Marco Rigobello and Roland Farrell for insightful comments on the manuscript, and Daniel Jaschke for support with software development. The tensor network simulations were performed with the Quantum Tea library [54, 55, 56]. The exact diagonalisation calculations were performed with QuSpin [57].